OpenStudy (anonymous):

Hey everybody im stuck on this problem Find the derivative of 8/(2+3e^(-x/2)) could someone please help?

OpenStudy (anonymous):

\[f(x)=\frac{8}{2+3e^{-x/2}}\] I don't know about you, but I prefer avoiding using the quotient rule whenever possible. To that end, I'll rewrite \[f(x)=8\left(2+3e^{-x/2}\right)^{-1}\] Now apply the chain rule to differentiate: \[f'(x)=-8\left(2+3e^{-x/2}\right)^{-2}\frac{d}{dx}\left[2+3e^{-x/2}\right]\] Think you can do the rest?

OpenStudy (anonymous):

What else do i need to do isnt that the derivative?

OpenStudy (anonymous):

There's still the matter of taking the remaining derivative. Have you learned the chain rule yet? \[f'(x)=-8\left(2+3e^{-x/2}\right)^{-2}\color{red}{\frac{d}{dx}\left[2+3e^{-x/2}\right]}\]